OpenStudy (anonymous):

The formula for degrees Celsius is: C = five over nine (F − 32), where F stands for degrees Fahrenheit. Which of the following equations shows the formula correctly solved for F?

OpenStudy (anonymous):

F = five over nine C + 32 F = nine over five C + 32 F = five over nine C − 32 F = nine over five C − 32